The angle of elevation of the top of cell tower B from the top of cell tower A is 30 degrees. The angle of depression of the foot of the cell tower B from the top of cell tower A is 60 degrees. The height of cell tower B is 100 meters. The foot of cell tower A and the foot of cell tower B is in the same horizontal plane.
What is the height of cell tower A?

Answer by Theo(13342)

the following picture shows the relationship between all the parts.

see below the picture for the analysis.

$$$

tan(30) = a/x

tan(60) = b/x

solve both these equations for x to get:

x = a/tan(30)

x = b/tan(60)

for ease of presentation, let T30 = tan(30) and T60 = tan(60).

because they both equal x, then they are both equal to each other so you get:

a/T30 = b/T60

multiply both sides of this equation by T30 to get:

a = T30/T60 * b

since a + b = 100, you can solve for b to get b = 100 - a.

substitute for b in the equation of a = T30/T60 * b to get:

a = T30/T60 * (100 - a)

simplify this to get:

a = (T30/T60 * 100) - (T30/T60 * a)

add (T30/T60 * a) to both sides of this equation to get:

a + (T30/T60 * a) = (T30/T60 * 100)

factor out the a on the left side of the equation to get:

a * (1 + T30/T60) = (T30/T60 * 100)

divide both sides of this equation by (1 + T30/T60) to get:

a = (T30/T60 * 100) / (1 + T30/T60)

since T30/T670 = 1/3, you can replace T30/T60 with 1/3 and solve for a to get:

a = (1/3 * 100) / (1 + 1/3) which becomes:

a = (1/3 * 100) / (4/3) which becomes:

a = (1/3) * (3/4) * 100 which becomes:

a = 25

since a + b = 100, this means that b = 75 which means that the height of tower A is equal to 75.